# The role of turbulence in setting the phase of the ISM and implications for the star formation rate

### v1

@@ Abstract
What regulates star formation in different regions of the Galaxy is still debated and especially the role of turbulence is not fully understood. In this work, we explore the link between star formation, turbulence and the thermal state of the multi-phase interstellar medium (ISM). We analyse a suite of stratified box simulations modelling a realistic ISM that aims to probe environments similar to those found in the Milky Way. Turbulence is injected through stellar feedback and an external large-scale driving force. We find that star formation can be either boosted or reduced when increasing the external driving strength, depending on the environment. When the density is sufficiently high, warm neutral gas naturally transitions to the cold phase, leading to high cold neutral medium (CNM) fractions of around 40%. Under these conditions, excessive large-scale driving leads to a slight reduction of the CNM fraction and an increase in the amount of gas that is thermally unstable. What limits the star formation in this regime is a reduced fraction of dense gas due to additional turbulent support against collapse. For low density regions, overdensities in which cooling is efficient are much rarer and we find that star formation is regulated by the formation of cold gas. In such cases, turbulence can significantly boost star formation by compressing gas in shocks and increasing the CNM fraction dramatically. In our simulations we see an increase from almost no CNM to up to a fraction of 15 % when including external turbulence driving; leading to an associated increase in the star formation rate. We provide a model to quantify this behaviour and predict the CNM fraction by combining the standard ISM cooling/heating model with the density PDF generated by turbulence. The change in the dominant limiting process for star formation between low- and intermediate-density environments provides a natural explanation for the observed break in the Kennicutt-Schmidt relation around column densities of 9 $\mathrm{M}_{\odot}$  pc-2.
@@ 1 Introduction
The relation becomes even more direct when considering only the molecular gas. Bigiel et al. (2008) have established a tight linear relation $\Sigma_{\mathrm{H}_{2}}\propto\Sigma_{\mathrm{SFR}}$, where H2 was estimated from CO emission (see also Pessa et al., 2021; Sun et al., 2023). In contrast, the relation between the SFR and the HI surface density is found to vary from galaxy to galaxy (Bigiel et al., 2008; Leroy et al., 2008; Shetty et al., 2014). Since only cold gas is relevant for the formation of stars (Krumholz et al., 2009; Klessen & Glover, 2016), regardless of whether it is in atomic or molecular form (Glover & Clark, 2012), we would not expect a universal relation of the SFR with the total amount of HI if the CNM fraction varies with environment.
@@
It is thought that the turbulence in molecular clouds is inherited from larger scales during their formation process (Heyer & Brunt, 2004; Brunt et al., 2009; Klessen & Hennebelle, 2010). What exactly drives the turbulence on these scales is still debated. In present-day spiral galaxies such as the Milky Way, numerical studies indicate that stellar feedback in the form of supernova (SN) possibly in combination with ionising radiation and winds is the dominant driving mechanism in most parts of the galaxy (Peters et al., 2017; Gatto et al., 2017; Kim et al., 2020; Rathjen et al., 2021). These studies reproduce levels of star formation in agreement with observations, without needing to invoke additional limiting processes. Recently, Colman et al. (2024) investigated simulated cloud properties in different environments and from different numerical setups. They found an overall trend of the turbulent velocity dispersion consistent with large-scale turbulent decay. Furthermore, at small spatial scales – and small molecular clouds – nearby SNe leave a noticeable imprint on the velocity dispersion.
@@
In this work, we study in more detail the effect of large-scale driving in Milky Way-like environments with column densities similar to what is found in the Solar neighbourhood. The paper is structured as follows. In Section 2, we describe the simulation setup. Our suite of eight simulations explores two values for the initial column density and various driving strengths. In Section 3, we analyse and compare the star formation activity in the different simulations. We find that, in some cases, driving can enhance star formation rather than reducing it. In Section 4, we investigate how turbulence affects the thermal balance in the simulations. Turbulence is found to influence the CNM and LNM fraction. This is then connected to the SFR in Section 5, where we take a closer look at the star formation relations. We identify two regimes, each with a different limiting process. In Section 6, we further discuss the generation of CNM through turbulence and construct the prototype of an analytical model to predict the CNM fraction. Finally, Section 7 summarises our findings and concludes the paper.
@@ 2 Simulations
As in previous work (Iffrig & Hennebelle, 2015; Colling et al., 2018; Brucy et al., 2020), we model a 1 kpc3 region of a galaxy with realistic ISM physics. The simulations used for this work are described in Colman et al. (2022). Additionally, we include four new runs with lower initial column density, to probe a wider variety of conditions.
2.1 Setup
We include an initial magnetic field with an orientation along the $x$-axis. The magnetic field strength has a Gaussian distribution in the vertical direction with the same thickness $z_{0}$ as the gas and a mid-plane value $B_{0}=7.62\,\mu$G, which is compatible with various estimates for the magnetic field in the Milky Way (Beck, 2003; Heiles & Crutcher, 2005; Heiles & Troland, 2005).
2.2 Numerics
2.3 Sink particles and stellar feedback
[…]
The massive stars are sources of stellar feedback. They emit ionising radiation from the position of the sink particle in which they formed and explode as a supernova at the end of their lifetime at a random location around the sink (Rosdahl et al., 2013; Geen et al., 2016; Iffrig & Hennebelle, 2017; Colling et al., 2018). The supernovae inject $4\times 10^{43}$ g cm s-1 momentum and $10^{51}$ erg of thermal energy, but the velocity and temperature of the surrounding gas are limited to 300 m s-1 and $10^{6}$ K respectively to prevent the time step from becoming very small. For low-mass stars, the UV output is negligible and their lifetimes are much longer than the simulation time. Due to their computational cost, we do not include stellar winds. Furthermore, similar numerical studies found that radiation is more important for the regulation of the SFR and that including winds further reduces the SFR by only a factor 2 (Rathjen et al., 2021).
2.4 Turbulence driving
Table 1: Overview of the simulations.
@@ 2.4 Turbulence driving
Different forcing strengths lead to different final global velocity dispersions. In Figure 2, we show the evolution of the velocity dispersion in the different simulations, measured in the disk. This quantity reflects the differences in energy injection rate from Figure 1.
From Figure 1 and 2 we can estimate the length of the initialisation phase and after how much time the initial conditions have been washed out. For the high-$\Sigma$ runs, the SN rate is converged after 50 Myr or less. The stronger the driving, the shorter the initialisation phase. In the low-$\Sigma$ runs, the evolution proceeds slower especially when there is little turbulence.
@@ 2.5 Cooling and heating model
The dominant heating process is photoelectric heating from small dust grains and PAHs, that is the emission of electrons from the grains after exposure to the interstellar radiation field. While UV radiation can free electrons in grains of all sizes, these electrons can escape more easily from small grains. Consequently, the photoelectric heating rate depends on the amount of small grains available and the strength of the far UV field $G_{0}$. Bakes & Tielens (1994) provide an analytical model as a function of $G_{0}$, the temperature $T$ and the electron density $n_{e}$, assuming a fixed grain size distribution. Even though the photoionisation of hydrogen is followed using Ramses-RT, the electron density used in the cooling model is calculated using the formula provided by Wolfire et al. (2003). Corrections for photoionisation in the WNM are made based on the work of Ferland (2003), as is done in Geen et al. (2016). A constant uniform UV background with a strength equal to 0.6 Habing units is assumed (Habing, 1969). In reality we expect the UV field to vary locally and to be proportional to the local SFR (Ostriker et al., 2010).
@@
Figure 4: The star formation history in each simulation. Top: evolution of the star formation rate. Bottom: evolution of star formation efficiency. Left: low-$\Sigma$. Right: high-$\Sigma$. The grey lines show the fits to estimate the global SFR of the simulation. The labels indicate the average SFR value found from these fits.
2.6 Simulation suite
The parameters of the different simulations are summarised in Table 1 and the final snapshot of each simulation is displayed in Figure 3. We can see a wide variety in terms of structure. The low-$\Sigma$ runs (top panels), especially the ones with no to weak driving, have a prominent diffuse gas component with a thin disk of denser material in which stars form. The high-$\Sigma$ runs (bottom panels) are more filamentary. Increasing the driving strength leads to more large-scale structure, seen as large overdensities and voids (Colman et al., 2022).
@@ 3.1 Star formation history
Using sink particles, we trace the star formation activity throughout the simulations. In the context of this work, sinks represent small star clusters rather than individual stars, and consists of a number of massive stars in addition to an unresolved collection of low-mass stars. Figure 4 displays the star formation history of each simulation. The top panels show the variation of the SFR in time, which exhibits large fluctuations. The bottom panel shows the evolution of the SFE, defined as the amount of gas which is in sinks compared to the total mass available at the beginning of the simulation. The general behaviour is different for the low-$\Sigma$ runs (10.2 $\mathrm{M}_{\odot}\mathrm{pc}^{-2}$, left panels) compared to the high-$\Sigma$ runs (19.1 $\mathrm{M}_{\odot}\mathrm{pc}^{-2}$, right panels) and we see two regimes emerge.
@@ 4.1 Gas phases
In the density PDF (top panel), the WNM and CNM phases can be identified as distinct bumps around a characteristic density of about 0.5 cm-3 and 30 cm-3 respectively. Note however that not all dense gas is cold. Dense shells around newly formed stars are heated and partially ionised by the UV radiation emitted by those stars. This makes photoelectric heating more efficient and brings the gas to WNM temperatures. Further investigation of this phenomenon requires more careful coupling between photoelectric heating, ionisation and the propagation of the local UV field in the simulations, which is beyond the scope of this work.
@@ 4.2 Fraction of gas in each phase
Figure 6: Evolution of the CNM and LNM mass fractions in full and empty circles respectively.
@@ 4.3 Dense gas
We further quantify the efficiency of the formation of dense gas in cold clouds in Figure 8, where we plot the ratio between the amount of CNM versus the amount of gas with densities greater than 100 cm-3. For a point in the bottom left part of the diagram, the amount of cold material is small and the conversion to high density gas inefficient. If a point is in he top part of the diagram, it means the conversion to dense gas is efficient. In the low-$\Sigma$ regime the conversion from cold to dense gas occurs at a roughly equal, albeit low, efficiency which is independent of the level of externally driven turbulence. While the amount of cold HI is similar for all high-$\Sigma$ runs, there is about a factor 2 to 3 less dense gas for the strong driving case. This brings the efficiency to the levels of the lower column density simulations. Strong driving hinders the creation of dense self-gravitating clumps. This is in line with predictions by analytical theories of gravo-turbulent fragmentation (Hennebelle & Chabrier, 2008; Hennebelle et al., 2024).
@@ 4.5 Caveats of the simulations
The transition point between WNM and CNM depends on the UV radiation field, metallicity and hydrogen ionisation rate (Wolfire et al., 2003). While our simulations do track the local UV photon field which is used to determine the ionisation of the gas, this information is not used in the cooling model where a uniform background UV field fixed to the solar neighbourhood value $G_{0}$ is assumed. Observations in the Milky-Way and nearby galaxies show that the UV field strength decreases as a function of galactic radius. The UV field in our low-$\Sigma$ runs may thus be too high compared to real low density regions in galaxies. A lower UV field implies less heating which would lead to naturally higher CNM fraction. Furthermore, in the majority of environments we expect the UV field to be directly coupled to the formation of new stars, especially massive ones. Test simulations where the uniform UV background is proportional to the SFR according to the recipe of Ostriker et al. (2010), indeed result in more structure formation and enhanced SFRs compared to the low-$\Sigma$ runs presented here. It is however unclear how local variations in the UV field may affect the overall CNM fraction. While we demonstrated a clear influence of turbulence under a fixed UV field, in practice regions with conditions similar to those in the low-$\Sigma$ runs may be rare.
Our simulations also do not take into account complex galactic effects such as differential rotation or the presence of spiral arms.

### v2

@@ Abstract
What regulates star formation in different regions of the Galaxy is still debated and especially the role of turbulence is not fully understood. In this work, we explore the link between star formation, turbulence and the thermal state of the multi-phase interstellar medium (ISM). We analyse a suite of stratified box simulations modelling a realistic ISM that aims to probe environments similar to those found in the Milky Way. Turbulence is injected through stellar feedback and an external large-scale driving force. We find that star formation can be either boosted or reduced when increasing the external driving strength, depending on the environment. When the density is sufficiently high or the initial UV background weak, warm neutral gas naturally transitions to the cold phase, leading to high cold neutral medium (CNM) fractions of around 30 – 40%. Under these conditions, excessive large-scale driving leads to a slight reduction of the CNM fraction and an increase in the amount of gas that is thermally unstable. What limits the star formation in this regime is a reduced fraction of dense gas due to additional turbulent support against collapse. For low density regions subject to significant external UV background, overdensities in which cooling is efficient are much rarer and we find that star formation is regulated by the formation of cold gas. In such cases, turbulence can significantly boost star formation by compressing gas in shocks and increasing the CNM fraction dramatically. In our simulations we see an increase from almost no CNM to up to a fraction of 15 % when including external turbulence driving; leading to an associated increase in the star formation rate. We provide a model to quantify this behaviour and predict the CNM fraction by combining the standard ISM cooling/heating model with the density PDF generated by turbulence. The change in the dominant limiting process for star formation between low-density /externally heated and intermediate-density /feedback heated environments could provides a natural explanation for the observed break in the Kennicutt-Schmidt relation around column densities of 9 $\mathrm{M}_{\odot}$  pc-2.
@@ 1 Introduction
The relation becomes even more direct when considering only the molecular gas. Bigiel et al. (2008) have established a tight linear relation $\Sigma_{\mathrm{H}_{2}}\propto\Sigma_{\mathrm{SFR}}$, where H2 was estimated from CO emission (see also Pessa et al., 2021; Sun et al., 2023). In contrast, the relation between the SFR and the HI surface density is found to vary from galaxy to galaxy (Bigiel et al., 2008; Leroy et al., 2008; Shetty et al., 2014). Since only cold gas is relevant for the formation of stars (Krumholz et al., 2009a; Klessen & Glover, 2016), regardless of whether it is in atomic or molecular form (Glover & Clark, 2012), we would not expect a universal relation of the SFR with the total amount of HI if the CNM fraction varies with environment.
@@
It is thought that the turbulence in molecular clouds is inherited from larger scales during their formation process (Heyer & Brunt, 2004; Brunt et al., 2009; Klessen & Hennebelle, 2010). What exactly drives the turbulence on these scales is still debated. In present-day spiral galaxies such as the Milky Way, numerical studies indicate that stellar feedback in the form of supernova (SN) possibly in combination with ionising radiation and winds is the dominant driving mechanism in most parts of the galaxy (Peters et al., 2017; Gatto et al., 2017; Colling et al., 2018; Rathjen et al., 2021). These studies reproduce levels of star formation in agreement with observations, without needing to invoke additional limiting processes. Recently, Colman et al. (2024) investigated simulated cloud properties in different environments and from different numerical setups. They found an overall trend of the turbulent velocity dispersion consistent with large-scale turbulent decay. Furthermore, at small spatial scales – and small molecular clouds – nearby SNe leave a noticeable imprint on the velocity dispersion.
@@
In this work, we study in more detail the effect of large-scale driving in Milky Way-like environments with column densities similar to what is found in the Solar neighbourhood. The paper is structured as follows. In Section 2, we describe the simulation setup. A first suite of eight simulations explores two values for the initial column density and various driving strengths. A constant UV background was used for these run. In Section 3, we analyse and compare the star formation activity in the different simulations in this first set. We find that, in some cases, driving can enhance star formation rather than reducing it. In Section 4, we investigate how turbulence affects the thermal balance in these simulations. Turbulence is found to influence the CNM and LNM fraction. In Section 5, we present a second set of simulations, now with a UV background which is time-dependent and proportional to the global SFR in the box. We repeat the previous analysis and stress the differences with the constant UV case. The connection to the SFR is explored in Section 6, where we take a closer look at the star formation relations. We identify two regimes, each with a different limiting process. In Section 7, we further discuss the generation of CNM through turbulence and construct the prototype of an analytical model to predict the CNM fraction. Finally, Section 8 summarises our findings and concludes the paper.
@@ 2 Simulations
Table 1: Overview of the simulations with constant UV background.
@@
As in previous work (Iffrig & Hennebelle, 2015; Colling et al., 2018; Brucy et al., 2020), we model a 1 kpc3 region of a galaxy with realistic ISM physics. The first set of simulations used for this work are described in Colman et al. (2022). Additionally, we extend this set to include four new runs with lower initial column density, to probe a wider variety of conditions. What is characteristic about these simulations is that they assume a constant UV background. To gain additional insight, we analysed a second set of simulations with a variable UV field. The details of the first set are described below. In Section 5.1, we discuss the differences with the second set.
2.1 Setup
[…]
2.2 Numerics
2.3 Sink particles and stellar feedback
When the gravitational collapse reaches the resolution limits, we introduce sink particles (Bate et al., 1995; Federrath et al., 2010b) according to the recipe prescribed by Bleuler & Teyssier (2014). After their birth, sinks accrete gas according to the threshold accretion scheme: gas which is above the sink formation threshold of $10^{4}$ cm-3 and within the accretion radius of 4 cells will be accreted. The high-$\Sigma$ simulations have been evolved for 60 Myr during which 2 to 5 $\%$ of the gas has been converted into stars. The low-$\Sigma$ runs are computationally less demanding and have varying final times, depending on the intensity of their star formation activity.
The massive stars are sources of stellar feedback. They emit ionising radiation from the position of the sink particle in which they formed and explode as a supernova at the end of their lifetime at a random location around the sink (Rosdahl et al., 2013; Geen et al., 2016; Iffrig & Hennebelle, 2017; Colling et al., 2018). The supernovae inject $4\times 10^{43}$ g cm s-1 momentum and $10^{51}$ erg of thermal energy, but the velocity and temperature of the surrounding gas are limited to 300 km s-1 and $10^{6}$ K respectively to prevent the time step from becoming very small. We have tested the impact of this procedure by comparing several statistics of runs with and without these restrictions and find limited differences. A comparison regarding the SFR has been presented in Brucy et al. (2023). In the Appendix B, we additionally show the velocity dispersion as a function of density. We find that the limitations do not significantly affect the results. The physical reason is that when very high velocities (typically larger than 300 km s-1) are needed, it is because the supernova explodes in a very tenuous medium. Such explosions have very limited impact on the dense gas, as studied for instance by Iffrig & Hennebelle (2015). For low-mass stars, the UV output is negligible and their lifetimes are much longer than the simulation time. Due to their computational cost, we do not include stellar winds. Furthermore, similar numerical studies found that radiation is more important for the regulation of the SFR and that including winds further reduces the SFR by only a factor 2 (Rathjen et al., 2021).
2.4 Turbulence driving
@@
Different forcing strengths lead to different final global velocity dispersions. In Figure 2, we show the evolution of the velocity dispersion in the different simulations, measured in the disk. This quantity reflects the differences in energy injection rate from Figure 1. We note that the values obtained in some of the low-$\Sigma$ runs are only around 2-3 km s-1, which is significantly lower than what is usually inferred in simulations and observations. This is due to the low SN rate in these runs. We will come back to this in Section 5.4. For the other simulations, our measured velocity dispersions are very similar to what is reported in other work, for instance Girichidis et al. (2016) and Kim et al. (2023) (their Fig. 19 and Table 5 respectively).
From Figures 1 and 2, we can estimate the length of the initialisation phase and after how much time the initial conditions have been washed out. For the high-$\Sigma$ runs, the SN rate is converged after 50 Myr or less. The stronger the driving, the shorter the initialisation phase. In the low-$\Sigma$ runs, the evolution proceeds slower especially when there is little turbulence.
@@ 2.5 Cooling and heating model
The dominant heating process is photoelectric heating from small dust grains and PAHs, that is the emission of electrons from the grains after exposure to the interstellar radiation field. While UV radiation can free electrons in grains of all sizes, these electrons can escape more easily from small grains. Consequently, the photoelectric heating rate depends on the amount of small grains available and the strength of the far UV field $G_{0}$. Bakes & Tielens (1994) provide an analytical model as a function of $G_{0}$, the temperature $T$ and the electron density $n_{e}$, assuming a fixed grain size distribution. Even though the photoionisation of hydrogen is followed using Ramses-RT, the electron density used in the cooling model is calculated using the formula provided by Wolfire et al. (2003). Corrections for photoionisation in the WNM are made based on the work of Ferland (2003), as is done in Geen et al. (2016). A constant uniform UV background with a strength equal to 0.6 Habing units is assumed (Habing, 1969). In reality, we expect the UV field to vary locally and to be proportional to the local SFR (Ostriker et al., 2010). The advantage of keeping the UV background constant and uniform is that it makes it easy to disentangle between the dynamical effects, induced by the stellar feedback and turbulent driving, and the thermal effect induced by UV heating. In Section 5, we will explore a second set of simulations where the UV heating is time-dependent and proportional to the star formation rate (but still uniform spatially).
@@
Figure 4: The star formation history in each simulation. Left: low-$\Sigma$. Right: high-$\Sigma$. Top: evolution of the star formation rate, smoothed over a time scale of 7.5 Myr for low-$\Sigma$ and 0.5 Myr for high-$\Sigma$. Bottom: evolution of star formation efficiency. The grey lines show the fits to estimate the global SFR of the simulation. The labels indicate the average SFR value found from these fits.
2.6 Simulation suite with constant UV field
The parameters of the first set of simulations are summarised in Tables 1 and the final snapshot of each simulation is displayed in Figure 3. We can see a wide variety in terms of structure. The low-$\Sigma$ runs (top panels), especially the ones with not too weak driving, have a prominent diffuse gas component with a thin disk of denser material in which stars form. The high-$\Sigma$ runs (bottom panels) are more filamentary. Increasing the driving strength leads to more large-scale structure, seen as large overdensities and voids (Colman et al., 2022).
@@ 3.1 Star formation history
Using sink particles, we trace the star formation activity throughout the simulations. In the context of this work, sinks represent small star clusters rather than individual stars, and consists of a number of massive stars in addition to an unresolved collection of low-mass stars. Figure 4 displays the star formation history of each simulation. The top panels show the variation of the SFR in time, which exhibits large fluctuations. The bottom panel shows the evolution of the SFE, defined as the amount of gas which is in sinks compared to the total mass available at the beginning of the simulation. The general behaviour is different for the low-$\Sigma$ runs (10.2 $\mathrm{M}_{\odot}\mathrm{pc}^{-2}$, left panels) compared to the high-$\Sigma$ runs (19.1 $\mathrm{M}_{\odot}\mathrm{pc}^{-2}$, right panels) and we see two regimes emerging.
@@ 4.1 Gas phases
In the density PDF (top panel), the WNM and CNM phases can be identified as distinct bumps around a characteristic density of about 0.5 cm-3 and 30 cm-3 respectively. Note however that not all dense gas is cold. Dense shells around newly formed stars are heated and partially ionised by the UV radiation emitted by those stars. This makes photoelectric heating more efficient and brings the gas to higher temperatures. Further investigation of this phenomenon requires more careful coupling between photoelectric heating, ionisation and the propagation of the local UV field in the simulations, which is beyond the scope of this work.
@@ 4.2 Fraction of gas in each phase
Figure 6: Evolution of the CNM and LNM mass fractions.
@@ 4.3 Dense gas
We further quantify the efficiency of the formation of dense gas in cold clouds in Figure 8, where we plot the ratio between the amount of CNM versus the amount of gas with densities greater than 100 cm-3. For a point in the bottom left part of the diagram, the amount of cold material is small and the conversion to high density gas inefficient. If a point is in he top part of the diagram, it means the conversion to dense gas is efficient. In the low-$\Sigma$ regime the conversion from cold to dense gas occurs at a roughly equal, albeit low, efficiency which is independent of the level of externally driven turbulence. While the amount of cold HI is similar for all high-$\Sigma$ runs, there is about a factor 2 to 3 less dense gas for the strong driving case. This brings the efficiency to the levels of the lower column density simulations. Strong driving hinders the creation of dense self-gravitating clumps. This is in line with predictions by analytical theories of gravo-turbulent fragmentation (Hennebelle et al., 2024).
@@ 4.5 Caveats of a uniform UV background
The transition point between WNM and CNM depends on the UV radiation field, metallicity and hydrogen ionisation rate (Wolfire et al., 2003). While our simulations do track the local UV photon field which is used to determine the ionisation of the gas, this information is not used in the cooling model where a uniform background UV field fixed to the solar neighbourhood value $G_{0}$ is assumed. Observations in the Milky-Way and nearby galaxies show that the UV field strength decreases as a function of galactic radius. The UV field in our low-$\Sigma$ runs may thus be too high compared to real low density regions in galaxies. A lower UV field implies less heating which would lead to naturally higher CNM fraction. Furthermore, in the majority of environments we expect the UV field to be directly coupled to the formation of new stars, especially massive ones. In the next section, we present the result from simulations where the uniform UV background is proportional to the SFR according to the recipe of Ostriker et al. (2010). We will see that this indeed result in more structure formation and enhanced SFRs compared to the low-$\Sigma$ runs presented in this section. It remains however unclear how local variations in the UV field may affect the overall CNM fraction. While we demonstrated a clear influence of turbulence under a fixed UV field, in practice regions with conditions similar to those in the low-$\Sigma$ runs may be rare.
Table 2: Overview of the simulations with time-dependent UV background.
Name $\Sigma$555initial gas column density in $\mathrm{M}_{\odot}$pc-2 $n_{0}$ 666initial mid-plane density in cm-3 $f_{\mathrm{rms}}$777normalisation factor for the turbulence driving strength $\sigma$ 888final velocity dispersion in km s-1
n0.66_rms00000 8.4 0.66 0 11
n0.66_rms00875 8.4 0.66 875 11
n0.66_rms01750 8.4 0.66 1750 11
n1_rms00000 12.7 1 0 13
n1_rms01750 12.7 1 1750 13
n1_rms03500 12.7 1 3500 14
n1_rms07000 12.7 1 7000 17
n1_rms14000 12.7 1 14000 20
Figure 9: The SFR in the simulations with variable UV background, smoothed over a time scale of 1 to 2 Myr.
Figure 10: Evolution of the CNM mass fractions in the second set of simulations.
Figure 11: Evolution of the mass-weighted 3D velocity dispersion in the second set of simulations, measured in a region between 100 pc above and 100 pc below the mid-plane.
5 Variable UV background
As a constant background UV field might not be sufficiently realistic, we now turn our attention to the second set of kpc-box simulations. An overview of their global properties is listed in Table 2.
5.1 Simulation suite with variables UV background
The second set of simulations uses different base parameters than the first set. Our goal in this section is to see if we recover similar behaviour as in the previous two sections but for different types of environments, rather than to directly compare the two simulation suites.
As in the first set, we explore two different initial column densities and various external large-scale turbulence driving strengths. The initial column densities are slightly lower, $\Sigma=8.4\,\mathrm{M}_{\odot}\ \mathrm{pc}^{-2}$ and $\Sigma=12.7\,\mathrm{M}_{\odot}\ \mathrm{pc}^{-2}$. These values are around the $\Sigma=10\,\mathrm{M}_{\odot}\ \mathrm{pc}^{-2}$ were we previously observed a boost in CNM and star formation with increasing external turbulence. The driving strengths have been adjusted accordingly (see Table 2). The initial magnetic field in the set is also lower, with a value of about 2 $\mu G$. The spatial resolution is also lower, with a maximum resolution of 1 pc. The sink formation threshold was adjusted accordingly.
A key difference is the UV background, which is variable in time and proportional to the SFR measured in the computational domain. Indeed, we expect the UV field to be proportional to the SFR (Ostriker et al., 2010). The mean FUV density relative to the solar neighbourhood value $G^{\prime}_{0}$ is
$G^{\prime}_{0}=\frac{\Sigma_{SFR}}{\Sigma_{SFR,\odot}}=\frac{\Sigma_{SFR}}{2.5\times 10^{-9}\mathrm{M}_{\odot}{\rm pc}^{-2}{\rm yr}^{-1}}.$ (1)
Additionally, we set a minimum UV field of 0.1 $G_{0}$. We use the implementation of Brucy et al. (2023).
These simulations track the formation of molecular hydrogen, but because it is not relevant for this study, we do not discuss it here.
5.2 Star formation rate
Figure 9 shows the SFR as a function of time for the eight time-dependent UV background simulations. The first thing to note is that the simulations without external turbulence driving now actively form stars. The SFR is much higher than the average value of $8\times 10^{-5}$ $\mathrm{M}_{\odot}$ yr-1 we found for the no-driving run with $\Sigma=$ 10.2 $\mathrm{M}_{\odot}$ pc-2 in Section 3. Also with driving, the values for the SFR seen in this second set of simulations are somewhat higher than those found in the low-$\Sigma$ runs with constant UV field and higher magnetic field.
In contrast to the results we found in Section 3, in this second set of simulations, the presence of turbulence driving always leads to a reduced SFR, even when the column density is low.
5.3 CNM fraction
The evolution of the CNM fraction in Figure 10 shows that it is generally higher than what we found for the low-$\Sigma$ runs in Section 4. The UV-field at the beginning of the simulation is set to a very low minimal value. This results in the rapid formation of CNM during the initialisation phase. When star formation commences after 20 to 30 Myr, the CNM fraction has reached values of around 30% and remains stable afterwards in most cases.
The presence of turbulence now clearly reduces the CNM fraction. For the simulations with $\Sigma=8.4\,\mathrm{M}_{\odot}\mathrm{pc}^{-2}$, the CNM fraction of the runs with external driving is only half of that for the no-driving case. Interestingly, the strength of the driving does not seem to play a role.
5.4 Velocity dispersion
Figure 11 displays the velocity dispersion as a function of time in the eight simulations of Table 2. A comparison with the runs presented in Figure 2 is particularly enlightening. Except for the two cases with the strongest driving, the velocity dispersions in this set of simulations reach about 11-13 km s-1, irrespectively of the forcing strength and initial column density. In contrast, when the UV background is kept constant, the runs with 10.2 M pc-2 present significantly lower velocity dispersion ranging from 2 km s-1 when no turbulent driving is applied, to 7 km s-1 when turbulent forcing operates. This stems from the much higher star formation rate that is achieved in the simulation with variable UV background. The stellar feedback, in particular the number of supernova explosions, is one to two orders of magnitude larger than in the case of constant UV background.
5.5 Constant or variable UV background?
The results found for simulations with a time-dependent UV background are somewhat different to those with the constant value presented in the previous sections. Here, the presence of external turbulence driving is always negative, both on the CNM fraction and SFR.
Although a UV background which is local and proportional to the SFR may be assumed to be the norm, there are situations where a constant UV field may be a good approximation. Let us remind that the mean free path of the UV photons in an ISM with a mean density of 0.3 cm-3 is about 1 kpc. This implies that for instance in the outskirts of galaxies, the ISM is heated by UV photons which have been emitted at several kpc distance. Accumulated over the galaxy and in a time-span considered by our simulations, the background can then be considered fairly constant.
Moreover, the choice of a constant UV background is a convenient one when comparing to analytical models. As will be addressed in Section 7, our numerical simulations are complemented by an analytical model of the warm and cold phase distribution for a specific UV field and ISM mean density. Simulations with a constant UV background provide important comparison cases.